# When Surfaces Lie: Exploiting Wrinkle-Induced Attention Shift to Attack Vision-Language Models

### 29 Mar 2026 v1

@@ 1. Introduction
Vision-Language Models (VLMs) have become a major paradigm for cross-modal understanding by jointly modeling visual and textual information. With the rapid progress of large-scale vision-language pre-training and multimodal instruction tuning, VLMs have shown strong generalization across zero-shot classification, image captioning, and visual question answering, and are increasingly viewed as a key route toward open-world visual understanding (Radford et al., 2021; Cherti et al., 2023; Li et al., 2022; Dai et al., 2023; Alayrac et al., 2022). Meanwhile, recent advances in test-time generalization, adversarial robustness, and multimodal safety suggest that VLMs are being deployed as increasingly general-purpose systems in complex environments (Cui et al., 2024; Lu et al., 2023; Zhang et al., 2022; Zanella and Ben Ayed, 2024). A careful understanding of their robustness under practically grounded distortions is therefore essential.
However, existing adversarial studies on VLMs still focus mainly on additive pixel-space perturbations, patch attacks, backdoor attacks, and joint image-text manipulations (Carlini and Wagner, 2017; Zhou et al., 2022; Zhang et al., 2025). Comparatively less attention has been paid to non-rigid structural perturbations that arise in realistic imaging conditions. In practice, inputs may be affected by paper creases, surface irregularities, local shadows, lighting changes, and perspective shifts, especially on approximately diffuse surfaces such as paper, posters, cloth, and recaptured screens. Prior work has shown that non-additive factors, including natural shadows, reflected light, illumination variation, physical-world transformations, and differentiable deformation, can also mislead visual models (Athalye et al., 2018; Zhong et al., 2022; Wang et al., 2023; Hsiao et al., 2024; Liu et al., 2025; Lu and others, 2024; Ruan et al., 2025; Engstrom et al., 2017). Compared with additive perturbations, structure-aware distortions may reorganize contours, textures, and spatial semantic cues in a qualitatively different way, thereby interfering with cross-modal alignment (Buzhinsky et al., 2023). This raises a central question: can wrinkle-like non-rigid structural distortions expose vulnerabilities of VLMs that are not fully revealed by conventional additive or illumination-dominated attacks?
@@
Our goal is to examine whether wrinkle-like structural distortion constitutes a distinct and practically meaningful robustness setting for multimodal systems, rather than merely introducing another perturbation type for VLMs. Such perturbations locally reorganize spatial layout rather than merely altering pixel values, and may therefore affect object boundaries, local semantic evidence, and image-text alignment simultaneously. We evaluate the proposed perturbation on zero-shot classification across multiple CLIP models, and then transfer the adversarial examples generated in this surrogate stage to image captioning and visual question answering. Through this unified pipeline, we systematically study the vulnerability of VLMs under wrinkle-like structural perturbations with mild surface-consistent appearance variation. The main contributions of this paper are as follows:
@@ 2.2. Structured and Physical Perturbations
Compared with pixel-level additive perturbations, structured and physically plausible adversarial disturbances better resemble practical imaging conditions and therefore have greater real-world relevance (Carlini and Wagner, 2017). Previous studies have shown that natural shadow, reflected light, illumination variation, and deformation can all mislead visual models (Zhong et al., 2022; Wang et al., 2023; Hsiao et al., 2024; Liu et al., 2025; Lu and others, 2024; Ruan et al., 2025). For approximately diffuse surfaces, classical reflectance analysis further suggests that changes in surface orientation may induce systematic luminance variation under illumination, providing a physical basis for mild appearance changes associated with wrinkle deformation (Basri and Jacobs, 2003). These findings indicate that effective real-world attacks need not appear as obvious noise, but may instead manifest as structural distortion, illumination change, or surface undulation. However, existing studies mainly emphasize shadow, lighting, or deformation-based perturbations, while wrinkle-like non-rigid distortions remain largely unexplored in current VLM robustness research (Zhong et al., 2022; Hsiao et al., 2024; Liu et al., 2025; Lu and others, 2024; Ruan et al., 2025).
@@ 3.2. Multi-Scale Wrinkle-Like Distortion
Perceptual similarity constraint. To balance attack effectiveness and visual plausibility, we introduce a perceptual similarity constraint to measure the visual closeness between the perturbed image and the original image. Since a single metric can hardly capture both structural preservation and high-level perceptual consistency, we combine SSIM (Wang et al., 2004) and LPIPS (Zhang et al., 2018) to construct a composite perceptual similarity metric:
(15) $S_{\mathrm{perc}}(x,x^{\prime})=0.7\cdot\bigl(1-S_{\mathrm{LPIPS}}(x,x^{\prime})\bigr)+0.3\cdot S_{\mathrm{SSIM}}(x,x^{\prime}).$
Here, $S_{\mathrm{SSIM}}$ reflects local structural preservation, while $S_{\mathrm{LPIPS}}$ measures perceptual distance in a high-level feature space. To make it consistent with SSIM in the optimization objective, we use $1-S_{\mathrm{LPIPS}}(x,x^{\prime})$ as a similarity-oriented term within $S_{\mathrm{perc}}$. Their combination jointly constrains structural preservation and perceptual fidelity, thereby encouraging visually plausible perturbations.
@@
(18) $F(\theta,\xi)=(1-\alpha)S_{\mathrm{ladv}}+\alpha S_{\mathrm{perc}},$
where $\alpha\in[0,1]$ balances attack effectiveness and perceptual quality. When the perturbation has already changed the model prediction, the fitness further favors candidates with higher perceptual quality among successful samples, and is defined as:
@@ 3.3. Genetic Algorithm Optimization
4. Experiments
4.1. Experimental Setup
Experimental Parameters. In our experiments, wrinkle-like perturbations are searched in a low-dimensional parameter space using a genetic algorithm under a black-box setting. The population size is set to 8, the maximum number of fitness evaluations for each image is set to 350 as the evaluation budget, and the perceptual weighting coefficient $\alpha$ in Eq. (18) is set to 0.3. The perturbation parameters include wrinkle-field generation parameters, displacement parameters, and related appearance parameters. Specifically, the strength ranges of the large- and medium-scale wrinkle components are set to $[0.4,0.8]$, while that of the small-scale component is set to $[0.3,0.5]$; the layer numbers of the large-, medium-, and small-scale components are set to $[2,4]$, $[4,6]$, and $[6,8]$, respectively. The decay and frequency parameters are searched within preset ranges for different scales. The displacement parameters $\gamma_{u}$ and $\gamma_{v}$ are both set to $[0.4,0.6]$, while the base brightness and brightness amplitude are set to $[0.4,0.8]$ and $[0.2,0.4]$, respectively. In the generative tasks, image captioning and visual question answering directly use the adversarial examples generated in the zero-shot classification stage for transfer evaluation. All experiments are conducted on an RTX 5090 GPU.
@@ 4.2. Comparative Experiments
As shown in Table 1, the proposed method achieves the strongest attack performance across all four CLIP models in terms of classification accuracy. Compared with clean samples, Ours (best) reduces the accuracy of OpenCLIP ViT-B/16, Meta-CLIP ViT-L/14, EVA-CLIP ViT-G/14, and OpenAI CLIP ViT-L/14 by 54, 40, 32, and 51 percentage points, respectively, consistently outperforming Natural Light Attack (Hsiao et al., 2024), Shadow Attack (Zhong et al., 2022), and ITA (Liu et al., 2025). Meanwhile, Ours (random) achieves the highest GPT-4o visual naturalness scores on all four models, indicating that the proposed perturbation can better preserve perceptual realism while maintaining substantial attack effectiveness. These results indicate that the proposed perturbation achieves a favorable trade-off between attack strength and visual naturalness.
@@ 5. Ablation Experiments
Effect of Multi-scale Wrinkle Components. We further analyze the effect of different wrinkle components on attack performance. As shown in Figure 5, the Full setting, which combines large-, medium-, and small-scale perturbations, consistently achieves the best or near-best ASR across all models. In contrast, Large only, Medium only, and Small only denote using only the corresponding single-scale component, while Large+Medium uses the large- and medium-scale components without the small-scale one. Among these reduced settings, Small only generally outperforms Large only and Medium only, suggesting that fine-scale perturbations play a more important role in disrupting model perception. However, Large+Medium remains inferior to Full, further highlighting the importance of the small-scale component. Overall, these results indicate that the proposed attack benefits from multi-scale complementarity rather than any single scale alone.
Figure 6. Ablation of the perceptual-similarity weighting coefficient $\alpha$ on EVA-CLIP ViT-G/14.
Effect of the Perceptual-Similarity Weighting Coefficient. We finally study the effect of the perceptual-similarity weighting coefficient $\alpha$ on attack performance and visual quality. As shown in Figure 6, increasing $\alpha$ generally improves ASR but leads to a noticeable decline in the GPT-4o score, indicating reduced visual quality. These results suggest that $\alpha$ governs the trade-off between attack effectiveness and naturalness. Considering both aspects, the choice of $\alpha=0.3$ is reasonable, as it provides a favorable balance between attack success and visual quality.
@@ 6. Conclusion
The robustness of vision-language models under wrinkle-like structural perturbations is systematically evaluated through a parametric attack framework based on non-rigid geometric deformation. Experimental results show consistent performance degradation in zero-shot classification, together with effective transfer to image captioning and visual question answering. Compared with Natural Light, Shadow, and ITA, the proposed method demonstrates stronger attack effectiveness across multiple models, while attention visualizations indicate that the perturbation shifts model focus away from semantically relevant regions. These findings suggest that vision-language models remain vulnerable to structured non-rigid perturbations beyond additive noise.

### 02 Apr 2026 v2

@@ Abstract.
$c$s: Computing methodologies Computer visionccs: Computing methodologies Image manipulationccs: Computing methodologies Machine learningccs: Information systems Multimedia content analysis
@@ 1. Introduction
Vision-Language Models (VLMs) have emerged as a major paradigm for cross-modal understanding by jointly modeling visual and textual information. Driven by advances in large-scale vision-language pre-training and multimodal instruction tuning, VLMs have demonstrated strong generalization across tasks such as zero-shot classification, image captioning, and visual question answering, and are increasingly regarded as a key route toward open-world visual understanding (Radford et al., 2021; Cherti et al., 2023; Li et al., 2022; Dai et al., 2023; Alayrac et al., 2022). Meanwhile, recent progress in test-time generalization, adversarial robustness, and multimodal safety suggests that VLMs are being deployed as increasingly general-purpose systems in complex environments (Cui et al., 2024; Lu et al., 2023; Zhang et al., 2022; Zanella and Ben Ayed, 2024). Consequently, understanding their robustness under practically grounded distortions has become increasingly important.
However, existing adversarial studies on VLMs still focus mainly on additive pixel-space perturbations, patch attacks, backdoor attacks, and joint image-text manipulations (Carlini and Wagner, 2017; Zhou et al., 2022; Zhang et al., 2025). Comparatively less attention has been paid to non-rigid structural perturbations that arise in realistic imaging conditions. In practice, inputs may be affected by paper creases, surface irregularities, local shadows, lighting changes, and perspective shifts, especially on approximately diffuse surfaces such as paper, posters, cloth, and recaptured screens. Prior work has shown that non-additive factors, including natural shadows, reflected light, illumination variation, physical-world transformations, and differentiable deformation, can also mislead visual models (Athalye et al., 2018; Zhong et al., 2022; Wang et al., 2023; Hsiao et al., 2024; Liu et al., 2025; Lu et al., 2024; Ruan et al., 2025; Engstrom et al., 2017). Compared with additive perturbations, structure-aware distortions may reorganize contours, textures, and spatial semantic cues in a qualitatively different way, thereby interfering with cross-modal alignment (Buzhinsky et al., 2023). For vision-language models, such distortions are especially important because they may simultaneously disturb local visual cues and image–text correspondence. Such effects may be amplified in multimodal systems, where errors in visual grounding can further propagate to downstream language outputs. This raises a central question: can wrinkle-like non-rigid structural distortions expose vulnerabilities of VLMs that are not fully revealed by conventional additive or illumination-dominated attacks?
@@
Our goal is to examine whether wrinkle-like structural distortion defines a distinct and practically meaningful robustness setting for multimodal systems, rather than simply introducing another perturbation type for VLMs. Unlike pixel-level perturbations, it locally reorganizes spatial layout and may therefore simultaneously affect object boundaries, local semantic cues, and image–text alignment. We evaluate the proposed perturbation on zero-shot classification across multiple CLIP models, and then transfer the generated adversarial examples to image captioning and visual question answering. Through this unified pipeline, we systematically investigate the vulnerability of VLMs under wrinkle-like structural perturbations with mild surface-consistent appearance variation. The main contributions of this paper are as follows:
@@ 2.2. Structured and Physical Perturbations
Compared with pixel-level additive perturbations, structured and physically plausible adversarial disturbances better resemble practical imaging conditions and therefore have greater real-world relevance (Carlini and Wagner, 2017). Previous studies have shown that natural shadow, reflected light, illumination variation, and deformation can all mislead visual models (Zhong et al., 2022; Wang et al., 2023; Hsiao et al., 2024; Liu et al., 2025; Lu et al., 2024; Ruan et al., 2025). For approximately diffuse surfaces, classical reflectance analysis further suggests that changes in surface orientation may induce systematic luminance variation under illumination, providing a physical basis for mild appearance changes associated with wrinkle deformation (Basri and Jacobs, 2003). These findings indicate that effective real-world attacks need not appear as obvious noise, but may instead manifest as structural distortion, illumination change, or surface undulation. However, existing studies mainly emphasize shadow, lighting, or deformation-based perturbations, while wrinkle-like non-rigid distortions remain largely unexplored in current VLM robustness research (Zhong et al., 2022; Hsiao et al., 2024; Liu et al., 2025; Lu et al., 2024; Ruan et al., 2025).
@@ 3.2. Multi-Scale Wrinkle-Like Distortion
Perceptual similarity constraint. To balance attack effectiveness and visual plausibility, we define a perceptual similarity constraint to measure the visual closeness between the perturbed image and the original image. Since a single metric cannot fully capture both structural preservation and high-level perceptual consistency, we combine SSIM (Wang et al., 2004) and LPIPS (Zhang et al., 2018) as
(15) $S_{\mathrm{perc}}(x,x^{\prime})=(1-\alpha_{1})\cdot S_{\mathrm{SSIM}}(x,x^{\prime})+\alpha_{1}\cdot\bigl(1-S_{\mathrm{LPIPS}}(x,x^{\prime})\bigr),$
where $\alpha_{1}\in[0,1]$ controls the trade-off between perceptual consistency and structural preservation. Here, $S_{\mathrm{SSIM}}$ measures local structural fidelity, while $1-S_{\mathrm{LPIPS}}(x,x^{\prime})$ serves as a similarity-oriented perceptual term.
@@
(18) $F(\theta,\xi)=(1-\alpha_{2})S_{\mathrm{ladv}}+\alpha_{2}S_{\mathrm{perc}},$
where $\alpha_{2}\in[0,1]$ balances attack effectiveness and perceptual quality. When the perturbation has already changed the model prediction, the fitness further favors candidates with higher perceptual quality among successful samples, and is defined as:
@@ 3.3. Genetic Algorithm Optimization
4. Experiments
4.1. Experimental Setup
Experimental Parameters. In our experiments, wrinkle-like perturbations are searched in a low-dimensional parameter space using a genetic algorithm under a black-box setting. The population size is set to 8, the maximum number of fitness evaluations for each image is set to 350 as the evaluation budget, the perceptual fusion coefficient $\alpha_{1}$ in the perceptual similarity term is set to 0.7, and the perceptual weighting coefficient $\alpha_{2}$ in Eq. (18) is set to 0.3. The perturbation parameters include wrinkle-field generation parameters, displacement parameters, and related appearance parameters. Specifically, the strength ranges of the large- and medium-scale wrinkle components are set to $[0.4,0.8]$, while that of the small-scale component is set to $[0.3,0.5]$; the layer numbers of the large-, medium-, and small-scale components are set to $[2,4]$, $[4,6]$, and $[6,8]$, respectively. The decay and frequency parameters are searched within preset ranges for different scales. The displacement parameters $\gamma_{u}$ and $\gamma_{v}$ are both set to $[0.4,0.6]$, while the base brightness and brightness amplitude are set to $[0.4,0.8]$ and $[0.2,0.4]$, respectively. In the generative tasks, image captioning and visual question answering directly use the adversarial examples generated in the zero-shot classification stage for transfer evaluation. All experiments are conducted on an RTX 5090 GPU.
@@ 4.2. Comparative Experiments
As shown in Table 1, the proposed method achieves the strongest attack performance across all four CLIP models in terms of classification accuracy. Compared with clean samples, Ours (best) reduces the accuracy of OpenCLIP ViT-B/16, Meta-CLIP ViT-L/14, EVA-CLIP ViT-G/14, and OpenAI CLIP ViT-L/14 by 54, 40, 32, and 51 percentage points, respectively, consistently outperforming Natural Light Attack (Hsiao et al., 2024), Shadow Attack (Zhong et al., 2022), and ITA (Liu et al., 2025). Meanwhile, Ours (random) achieves the highest GPT-4o visual naturalness scores on all four models, indicating that the proposed perturbation can better preserve perceptual realism while maintaining substantial attack effectiveness. Notably, this advantage is not limited to a single backbone, but remains consistent across CLIP variants with different architectures and scales. This suggests that the proposed perturbation captures a transferable structural vulnerability in visual encoding rather than exploiting a model-specific weakness. These results indicate that the proposed perturbation achieves a favorable trade-off between attack strength and visual naturalness.
@@ 5. Ablation Experiments
Effect of Multi-scale Wrinkle Components. We further analyze the effect of different wrinkle components on attack performance. Here, Full combines large-, medium-, and small-scale perturbations; Large only, Medium only, and Small only use only the corresponding single-scale component; and Large+Medium uses the large- and medium-scale components without the small-scale one. As shown in Figure 5, Full achieves the best or near-best ASR across all models. Among the reduced settings, Small only generally outperforms Large only and Medium only, while Large+Medium still remains inferior to Full. These results indicate that the proposed attack benefits from multi-scale complementarity, with the small-scale component playing an important role.
Effect of the Perceptual-Fusion Coefficient. Using EVA-CLIP ViT-G/14 as an example, we further study the effect of the perceptual-fusion coefficient $\alpha_{1}$. As shown in Figure 6 (a), changing $\alpha_{1}$ causes limited variation in ASR, while the GPT-4o score changes more noticeably. In particular, $\alpha_{1}=0.7$ achieves the best visual quality while maintaining competitive attack performance. This indicates that assigning a higher weight to the LPIPS-based perceptual term yields a better balance between structural preservation and high-level perceptual consistency. Therefore, $\alpha_{1}=0.7$ is a reasonable choice.
Figure 6. Ablation study of the weighting coefficients $\alpha_{1}$ and $\alpha_{2}$ on EVA-CLIP ViT-G/14.
Effect of the Attack–Perceptual Weighting Coefficient. Using EVA-CLIP ViT-G/14 as an example, we study the effect of the attack–perceptual weighting coefficient $\alpha_{2}$ on attack performance and visual quality. As shown in Figure 6 (b), a larger $\alpha_{2}$ generally improves ASR, while causing a clear drop in the GPT-4o score, indicating lower visual quality. This suggests that increasing the weight of the perceptual term improves attack effectiveness, but weakens the visual naturalness of the perturbations. Considering this trade-off, $\alpha_{2}=0.3$ is a reasonable setting, as it provides a balanced compromise between attack success and visual quality.
@@ 6. Conclusion
The robustness of vision-language models under wrinkle-like structural perturbations is systematically evaluated through a parametric attack framework based on non-rigid geometric deformation. Experimental results show consistent performance degradation in zero-shot classification, together with effective transfer to image captioning and visual question answering. Compared with the baseline methods, the proposed method demonstrates stronger attack effectiveness across multiple models, while attention visualizations indicate that the perturbation shifts model focus away from semantically relevant regions. These findings suggest that vision-language models remain vulnerable to structured non-rigid perturbations beyond additive noise.